### v3

# Very Basic Set theory

Very Basic Set theory
Doeko H. Homan
(Date: August 18, 2024)
George Boolos describes the naive conception of set: ”Any predicate has an extension” (Boolos [1971] pp.215-231). Then Russell’s paradox is used to prove the inconsistency of naive set theory. Russell’s paradox reads:
”No set can contain all and only those sets which do not contain themselves”.
We call a set which contains itself an ‘individual’.
In this article we show Russell’s paradox is logically equivalent to:
For every set $s$, an individual is contained in $s$ or a set that is not an individual
is not contained in $s$.
Therefore Very basic set theory is a set theory with individuals.
1. What are individuals?
In (Kanamori [2003] pp.285-288) Akihiro Kanamori describes that with the ‘empty
extension’ and the ‘axiom of restriction’ there is no need for individuals (also
known as atoms or urelemente). Individuals ”are objects distinct from the
null set yet having no members and capable of belonging to sets”.
In (Boolos [1971] pp.216-217) ”individuals (= non-sets)”.
In section 4.6.5 (Mendelson [2015]) Elliott Mendelson explains:
”The words ‘individual’ and ‘atom’ are sometimes used as synonyms for
‘urelement’. Thus, urelements have no members”. And in a footnote (p.304):
”Zermelo’s 1908 axiomatization permitted urelements”.
In (Zermelo [1908] pp.262-265) Ernst Zermelo describes in ”10. Theorem.”:
”Für jedes Element $x$ von $M$ ist es definit, ob $x$ $\varepsilon$ $x$ oder nicht; diese Möglichkeit
$x$ $\varepsilon$ $x$ ist an und für sich durch unsere Axiome nicht ausgeschlossen.”
Thus ‘$x$ $\varepsilon$ $x$ or not’ is definite; the possibility ‘$x$ $\varepsilon$ $x$’ is not excluded by our axioms.
Therefore Zermelo’s 1908 axiomatization permitted sets which contain themselves.
Sets are denoted by $s$, $t$, $u$, …. Symbol $\in$ is the ‘membership relation’ symbol. Then $s\in t$ means $t$ contains $s$ or $s$ is a member of $t$. The negation of $s\in t$ is denoted by $s\notin t$. The membership relation is definite, that is for every $s$ and for every $t$ ($s\in t$ or $s\notin t$) but never ($s\in t$ and $s\notin t$). If $s\in s$ then $s$ is an individual.
We prove Russell’s paradox.
Assume exists $s$ such that for every $u$ ($u\in s$ if and only if $u\notin u$).
The formula applies to every $u$ thus we can choose $s$ therefore
exists $s$ such that ($s\in s$ if and only if $s\notin s$).
The question ‘is $s$ contained in $s$’ or ‘is $s$ not contained in $s$’ is definite but there is
no answer: If $s\in s$ then $s\notin s$ and vice versa. Thus $s$ is nonexisting. Therefore:
”No set can contain all and only those sets which do not contain themselves”. $\Box$
The membership relation is definite thus for every $s$ ($s\in s$ or $s\notin s$).
If $s\in s$ then exists $u$ such that ($u\in s$ and $u\in u$),
if $s\notin s$ then exists $u$ such that ($u\notin s$ and $u\notin u$) therefore:
For every $s$ exists $u$ such that (($u\in s$ and $u\in u$) or ($u\notin s$ and $u\notin u$)).
That is logically equivalent to Russell’s paradox:
It is not the case that exists $s$ such that for every $u$ ($u\in s$ if and only if $u\notin u$).
Thus for every $s$, an individual is contained in $s$ or a set that is not an individual
is not contained in $s$. $\Box$
Zermelo’s 1908 axiomatization permitted individuals. And the prove of Russell’s paradox includes individuals and other sets. But a set which contains itself, an individual, is problematic. (Boolos [1971] p.217, pp.219-220):
”perhaps the mind ought to boggle at the idea of something’s containing itself.”
”It is important to realize how odd the idea of something’s containing itself is.”
”Here are some things. Now we bind them up into a whole. Now we have a set.”
However, ‘binding up some things’ or ‘containing something’ is not the definition of a set. (Boolos [1971] p.215):
”Cantor also defined a set as a ‘many, which can be thought of as one, i.e. a
totality of definite elements that can be combined into a whole by a law.’ ”
Then ”which can be thought of as one,” a set is a ‘mental picture’ in your mind. Therefore we prefer to write ‘constituent of a set’ rather than ‘contained in a set’.
Set theory is a language that follows the rules of logic to communicate the mental picture of a set with the physical world. Logic treats every possibility. Individuals are possible. Then, hopefully, someone reading or listening that language forms a correct mental picture in her/his mind. That is called ‘experience’.
What is an individual? ‘Set $s$ is an individual’ is what is called in (Wittgenstein [1999]) ”an atomic fact”. In the Introduction (p.12) Bertrand Russell explains:
”If an atomic fact is analyzed as fully as possible (theoretical, not practical
possibility is meant) the constituents finally reached may be called simples
or objects.”
”(…) the naming of simples is shown to be what is logically first in logic.”
Therefore in Very basic set theory the naming of individuals is logically first.
2. What is a set?
At an early age you develop the mental picture of ‘what is a set’. You belong to a family or a clan, to the inhabitants of a village or a particular region. Then experience shows there are things that constitute a set. The members of a set are the constituents of that set. There is a membership relation between sets.
Very basic set theory is based on
the world of individuals,
the membership relation is definite,
the basic properties of a set are the constituents of that set and the sets that set is a member of,
the only constituent of an individual is that individual itself.
Experience shows a set ‘is equal to’ or ‘is not equal to’ another set. Sets with different constituents are different sets, and different sets have different constituents.
3. Very basic set theory
A ‘formula’ is a description of properties of sets. Thus $s\in t$ is a formula. Formulas are denoted by $\Phi$, $\Psi$, $\ldots$ and are built up from formulas by means of negation $\neg\Phi$ (‘not $\Phi$’) and conjunction $\Phi\wedge\Psi$ (‘$\Phi$ and $\Psi$’). Negation has higher precedence than conjunction. The parenthesis (((( and )))) preclude ambiguity and improve legibility. Then $s\notin t$ is short for the formula $\neg(s\in t)$.
If $\Phi$ is a formula then $(\Phi)$, $\forall s(\Phi)$ and $\exists s(\Phi)$ are formulas. $\forall s(\Phi)$ means ‘for every set $s$ $\Phi$’, and $\exists s(\Phi)$ means ‘exists set $s$ such that $\Phi$’. Then e.g. $\neg\exists s\forall u(\Phi)$ is short for the formula $\neg(\exists s(\forall u(\Phi)))$.
If there is $s$ in $\Phi$ not guided in $\Phi$ by $\forall s$ or by $\exists s$ then ‘$s$ is free in $\Phi$’. If $s$ is free in $\Phi$ and $\neg\exists s(\Phi\wedge\neg\Phi)$ then ‘$\Phi$ is definite’.
We define the logical connectives $\vee$, $\rightarrow$ and $\leftrightarrow$. The precedence of each of it is lower than conjunction.
$\Phi\vee\Psi$ (‘$\Phi$ or $\Psi$ or both’) is short for the formula ¬($\neg(\neg\Phi\wedge\neg\Psi)$,
$\Phi\rightarrow\Psi$ (‘$\Phi$ implies $\Psi$’) is short for the formula ($\neg\Phi\vee\Psi$),
$\Phi\leftrightarrow\Psi$ (‘$\Phi$ if and only if $\Psi$’) is short for the formula ($(\Phi\rightarrow\Psi)\wedge(\Psi\rightarrow\Phi))$.
If $\Phi$ and $\Psi$ are definite then $\neg\Phi$, $\Phi\wedge\Psi$, $\Phi\vee\Psi$, $\Phi\rightarrow\Psi$ and $\Phi\leftrightarrow\Psi$ are definite.
Then Russell’s paradox is the formula $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$, and is logically equivalent to $\forall s\exists u((u\in s\wedge u\in u)\vee(u\notin s\wedge u\notin u))$.
The naive conception of set is inconsistent: formula $\exists s\forall u(u\in s\leftrightarrow u\notin u)$ is not definite. Therefore we postulate ‘specification’ (also known as ‘separation’).
  • if $u$ is free in $\Phi$ and $\Phi$ is definite then $(u\in s\wedge\Phi)$ is definite and
    (specification) $\forall s(\exists u(u\in s\wedge\Phi)\rightarrow\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge\Phi)))$.
Thus the conjunction of mutually exclusive properties does not constitute a set.
We prove: If $\exists u(u\in s\wedge u\in u)$ then exists $v$ whose members are all and only
those individuals that are constituent of $s$.
Formula $(u\in s\wedge u\in u)$ is definite and is not the conjunction of mutually exclusive properties. Apply ‘specification’ to $s$ to find $v$
$\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge u\in u))$. Then an individual is a constituent of $v$.
The formula applies to every $u$ thus we can choose $v$
$\exists v(v\in v\leftrightarrow(v\in s\wedge v\in v))$ therefore
$\exists v((v\notin v\vee(v\in s\wedge v\in v))\wedge((v\notin s\vee v\notin v)\vee v\in v))$.
Then $\exists v(v\notin v\vee v\in s)$. Thus $v$ is not an individual, or $v$ is an individual and a constituent of $s$. A set of all individuals is not excluded. $\Box$
We prove ‘the set of all sets’ is nonexisting.
If $\exists u(u\in s\wedge u\notin u)$ then formula $(u\in s\wedge u\notin u)$ is definite and is not the conjuction of mutually exclusive properties.
Apply ‘specification’ to $s$ to find $v$
$\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge u\notin u))$. Then a set is a member of $v$.
The formula applies to every $u$ thus we can choose $v$
$\exists v(v\in v\leftrightarrow(v\in s\wedge v\notin v))$ therefore
$\exists v((v\notin v\vee(v\in s\wedge v\notin v))\wedge((v\notin s\vee v\in v)\vee v\in v))$.
Then $\exists v(v\notin v\wedge v\notin s)$. Thus $v$ is not an individual and $v$ is not a member of $s$. Therefore ‘the set of all sets’ is nonexisting. $\Box$
4. More Axioms
Experience shows a set is equal to or a set is not equal to another set.
We define equality for sets.
Definition 1.
Set $s$ ‘is equal to’ set $t$, denoted by $s=t$ if $\forall u(u\in s\leftrightarrow u\in t)$.
Set $s$ ‘is not equal to’ set $t$, denoted by $s\neq t$ if $\neg(s=t)$. $\Box$
The axioms for equality are
  • (equality) $\forall s\forall t(s=t\rightarrow\forall u(s\in u\leftrightarrow t\in u))$,
  • (individuals) $\forall s(s\in s\rightarrow\forall u(u\neq s\rightarrow u\notin s))$.
Thus if $s=t$ then $s$ and $t$ have the same constituents and $s$ and $t$ are members of the same sets. And the only constituent of an individual is that individual itself.
Then $\forall s\forall t((s\in s\wedge t\notin t)\rightarrow(s\neq t\wedge t\notin s))$.
Any pair of sets constitute a set. We postulate ‘pairs’.
  • (pairs) $\forall s\forall t\exists v\forall u(u\in v\leftrightarrow(u=s\vee u=t))$.
Set $v$ is denoted by $\{s,t\}$ thus $\{s,t\}=\{t,s\}$. The ‘singleton’ $\{s\}$ of $s$ is $\{s,s\}$.
$s\notin s\rightarrow\{s\}\neq s$ otherwise $s\in s$. Therefore $s\notin s\rightarrow\{s\}\notin\{s\}$,
$s\in s\rightarrow\{s\}=s$, the singleton of an individual is that individual itself,
$s\neq t\rightarrow\{s,t\}\notin\{s,t\}$. Thus a pair of different sets is not an individual.
However, if $s\neq t$ and $(s\in t\wedge t\in s)$ then $s\notin s$ and $t\notin t$. Then exists $\{s,t\}$ and there is the ‘vicious membership circle’ $(s\in t\wedge t\in s)$.
The naming of individuals is logically first. Thus every set has a constituent, and vicious membership circles do not exist. Therefore we postulate ‘regularity’.
  • (regularity) $\forall s\exists v(v\in s\wedge(v\notin v\rightarrow\forall u((u\in s\wedge u\in v)\rightarrow u\in u)))$.
If $s$ is an individual or a set of individuals then $v$ is an individual. Otherwise exists $v$ and $v$ is not an individual and $s$ and $v$ only have individuals in common.
And it is not the case that exists $s$ such that
$\forall v(v\in s\wedge v\notin v\rightarrow\exists u(u\in s\wedge u\in v\wedge u\notin u))$.
Therefore any sequence of sets $s$, $t$, $u$, $v$, … such that each term is a member of the previous one $t\in s$, $u\in t$, $v\in u$ is nonexisting. Thus $(t\in s\wedge s\in t)$ is nonexisting therefore $\forall s\forall t(s\neq t\rightarrow(s\notin t\vee t\notin s))$.
The members of the members of a set constitute a set. We postulate ‘union’.
  • (union) $\forall s\exists v\forall u(u\in v\leftrightarrow\exists t(t\in s\wedge u\in t))$.
Set $v$ is denoted by $\bigcup s$. If $s$ is an individual or a set of individuals then $\bigcup s=s$. The ‘union $s\cup t$’ is defined by $s\cup t=\bigcup\{s,t\}$. Then $\bigcup\{\{s\},\{t\}\}=\{s,t\}$ and $s\cup(t\cup u)=(s\cup t)\cup u$. If $s$ and $t$ are individuals then $s\cup t=\{s,t\}$.
An ‘infinite’ sequence of different sets $s$, $\bigcup s$, $\bigcup(\bigcup s)$, … is nonexisting.
Any pair of different consecutive sets $\{w,\bigcup w\}$ implies $\exists v(v\in w\wedge v\notin v)$ otherwise $\bigcup w=w$. If $z=s\cup\bigcup s\cup\bigcup(\bigcup s)\ldots$ then
$\forall w((w\in z\wedge w\notin w)\rightarrow\exists v(v\in w\wedge v\in\bigcup z\wedge v\notin v))$.
Then $\forall v(v\in\bigcup z\rightarrow v\in z)$ thus $z$ contradicts regularity.
Therefore the sequence terminates: $\exists w(\bigcup w=w)$. Set $z$ is called ‘the transitive closure’ of $s$. Thus for every set the transitive closure exists. $\Box$
5. Conclusion
Very basic set theory is a set theory with individuals
the only constituent of an individual is that individual itself,
Russell’s paradox is logically equivalent to a formula with individuals,
the conjunction of mutually exclusive properties does not constitute a set,
the naming of individuals is logically first,
for every set the transitive closure exists.
Compared with ‘mainstream set theory’ a big difference is ‘what is logically first’. The logic of set theory allows for individuals, and Russell’s paradox is logically equivalent to a formula with individuals. Therefore we choose for ‘the naming of individuals is logically first’, not for the postulate of an ‘empty set’.
There are more differences. The naive conception of set is inconsistent thus axiom ‘specification’ is modified. The axiom ‘powerset’ is as usual but Cantor’s Theorem does not hold.
  • (powerset) $\forall s\exists v\forall u(u\in v\leftrightarrow\forall t(t\in u\rightarrow t\in s))$.
Set $v$ is denoted by ${\mathcal{P}}s$. Then ${\mathcal{P}}\{s\}=\{\{s\}\}$ thus the ‘number of members’ of $\{s\}$ is equal to the ‘number of members’ of ${\mathcal{P}}\{s\}$.
Cantor’s Theorem (Zermelo [1908] p.276) reads: For every $s$,
the ‘number of members’ of $s$ is less than the ‘number of members’ of ${\mathcal{P}}s$.
Thus Cantor’s Theorem does not hold.
The definition of equality and axiom ‘pairs’ are only meaningful in combination with sets that are logically first. Thus with axiom ‘regularity’. And ‘regularity’ is used with ‘union’ to prove that for every set the transitive closure exists.
In the last section we define ‘number sequences’ satisfying the Peano Axioms (PA) for natural numbers. The first axiom of PA reads: There is a ‘first natural number’. Any pair of individuals is a first number of a natural number sequence. Therefore ‘the set of all natural numbers’ is nonexisting. However, a set of all natural numbers with a specified first number is not excluded.
If in the world only one pair of individuals exists then indeed a set of all natural numbers with that pair of individuals as first number can postulated to exist. That is the way of ‘mainstream set theory’: There is no need for individuals, and the ‘empty set’ is postulated as ‘first natural number’.
6. Natural numbers
Definition 2.
Set $s$ is a ‘subset’ of $t$, denoted by $s\subseteq t$, if $\forall u(u\in s\rightarrow u\in t)$.
$s$ is ‘transitive’ if $\forall u(u\in s\rightarrow u\subseteq s)$, or equivalently $\bigcup s\subseteq s$. $\Box$
Every individual and every set of individuals is transitive. If $s$ is transitive then $\bigcup s$ and $s\cup\{s\}$ are transitive.
Theorem 1.
If $s$ is transitive and $\exists u(u\in s\wedge u\notin u)$ then a set of individuals is a
member of $s$: $\exists v(v\in s\wedge v\notin v\wedge\forall u(u\in v\rightarrow u\in u))$.
Proof. Apply regularity to $s$ to find $v\in s\wedge v\notin v$ such that
$\forall u((u\in s\wedge u\in v)\rightarrow u\in u)$. Set $s$ is transitive and $v\in s$ thus $\forall u(u\in v\rightarrow u\in s)$.
Therefore $\exists v(v\in s\wedge v\notin v\wedge\forall u(u\in v\rightarrow u\in u))$. $\Box$
If $u$ and $v$ are the only individuals belonging to transitive set $s$ then $\{u,v\}$ is the only possible set of individuals belonging to $s$. Therefore $\{u,v\}\in s\cup\{s\}$.
If $s$ and $t$ are transitive sets with transitive members and the only individuals belonging to $s$ or to $t$ are $u$ and $v$, thus $\forall w((w\in s\cup t\wedge w\in w)\rightarrow w\in\{u,v\})$, then the ‘law of trichotomy’ for $s$ and $t$ is easily proved and reads
$(s\notin s\wedge t\notin t)\rightarrow(s\in t\vee s=t\vee t\in s).$
If $s$ is a transitive set with transitive members then $\bigcup s$ is a transitive set with transitive members and $s\notin s\rightarrow\bigcup s\notin\bigcup s$. And $s\notin\bigcup s$ otherwise $s\in s$. Therefore $\bigcup s=s$ or $s=\bigcup s\cup\{\bigcup s\}$.
It is straightforward to define a number sequence satisfying the Peano Axioms for natural numbers. Natural numbers are also known as ‘finite ordinals’.
Definition 3.
Set $s$ is a ‘natural number with first number $\alpha$’ if $s$ is a transitive
set with transitive members, $\alpha$ is a pair of individuals and
$\bigcup\alpha=\alpha\wedge s\notin\alpha\wedge\forall u(u\in s\cup\{s\}\rightarrow(u\in\alpha\cup\{\alpha\}\vee\bigcup u\in u)).$
The ‘successor’ of natural number $s$ is $s\cup\{s\}$. $\Box$
If $z$ (‘zero’), $x$ and $y$ are individuals then there is number sequence $S$ with first number (in physical reality called a ‘unit of measurement’) $\{z,x\}$ (‘x coordinate’) and there is number sequence $T$ with first number (unit of measurement) $\{z,y\}$ (‘y coordinate’). All elements of both sequences have $z$ in common.
If $u$ belongs to the $S$ sequence or to the $T$ sequence then $x\in u$ implies $u$ is an $x$ coordinate and $y\in u$ implies $u$ is an $y$ coordinate.