### v4

# Very Basic Set theory

Very Basic Set theory
Doeko H. Homan
(Date: January 15, 2025)
George Boolos describes the naive conception of set ”Any predicate has an extension” (Boolos [1971] pp.215-231). Then Russell’s paradox is used to prove the inconsistency of naive set theory. Russell’s paradox reads
”No set can contain all and only those sets which do not contain themselves”.
We call a set which contains itself an ‘individual’.
In this article we show Russell’s paradox is logically equivalent to:
For every set $s$, an individual is contained in $s$ or a set that is not an individual
is not contained in $s$.
Therefore Very basic set theory is a set theory with individuals.
1. What are individuals?
In (Kanamori [2003] pp.285-288) Akihiro Kanamori describes that with the ‘empty
extension’ and the ‘axiom of restriction’ there is no need for individuals (also
known as atoms or urelemente). Individuals ”are objects distinct from the
null set yet having no members and capable of belonging to sets”.
In (Boolos [1971] pp.216-217) ”individuals (= non-sets)”.
In section 4.6.5 (Mendelson [2015]) Elliott Mendelson explains
”The words ‘individual’ and ‘atom’ are sometimes used as synonyms for
‘urelement’. Thus, urelements have no members”. And in a footnote (p.304)
”Zermelo’s 1908 axiomatization permitted urelements”.
In (Zermelo [1908] pp.262-265) Ernst Zermelo describes in ”10. Theorem.”
”Für jedes Element $x$ von $M$ ist es definit, ob $x$ $\varepsilon$ $x$ oder nicht; diese Möglichkeit
$x$ $\varepsilon$ $x$ ist an und für sich durch unsere Axiome nicht ausgeschlossen.”
Thus ‘$x$ $\varepsilon$ $x$ or not’ is definite; the possibility ‘$x$ $\varepsilon$ $x$’ is not excluded by our axioms.
Therefore Zermelo’s 1908 axiomatization permitted sets which contain themselves.
Sets are denoted by $s$, $t$, $u$, …. Symbol $\in$ is the ‘membership relation’ symbol. Then $s\in t$ means $s$ is ‘contained’ in $t$. The negation of $s\in t$ is denoted by $s\notin t$. The membership relation is definite, that is for every $s$ and for every $t$ either $s\in t$ or $s\notin t$ but not both. If $s\in s$ then $s$ is an individual.
We prove Russell’s paradox
”No set can contain all and only those sets which do not contain themselves”.
Assume exists $s$ such that for every $u$ ($u\in s$ if and only if $u\notin u$).
The formula applies to every $u$ thus we can choose $s$ therefore
exists $s$ such that ($s\in s$ if and only if $s\notin s$).
The question ‘is $s$ contained in $s$’ or ‘is $s$ not contained in $s$’ is definite but there is
no answer: If $s\in s$ then $s\notin s$ and vice versa. Therefore $s$ is nonexisting. $\Box$
For every $s$ and for every $t$ the membership relation $s\in t$ is definite. Thus
if $s\in s$ then exists $u$ such that ($u\in s$ and $u\in u$),
if $s\notin s$ then exists $u$ such that ($u\notin s$ and $u\notin u$). Then for every $s$, an individual is contained in $s$ or a set that is not an individual is not contained in $s$.
Zermelo’s 1908 axiomatization permitted individuals. And the prove of Russell’s paradox includes individuals and other sets. But a set which contains itself, an individual, is problematic. (Boolos [1971] p.217, pp.219-220)
”perhaps the mind ought to boggle at the idea of something’s containing itself.”
”It is important to realize how odd the idea of something’s containing itself is.”
”Here are some things. Now we bind them up into a whole. Now we have a set.”
However, ‘binding up some things’ or ‘containing something’ is not the definition of a set. (Boolos [1971] p.215)
”Cantor also defined a set as a ‘many, which can be thought of as one, i.e. a
totality of definite elements that can be combined into a whole by a law.’ ”
Then ”which can be thought of as one,” thus a set is a ‘mental picture’ in your mind. Therefore we prefer to write ‘constituent of a set’, not ‘contained in a set’.
Set theory is a language that follows rules of logic to communicate the mental picture of a set with the physical world. Logic treats every possibility, and $s\in s$ is possible. Then someone reading or listening that language forms a mental picture of a set in her/his mind. The picture in her/his mind agrees or disagrees with the picture in your mind. We call that ‘experience’.
What is an individual? ‘Set $s$ is an individual’ is what is called in (Wittgenstein [1999]) ”an atomic fact”. In the Introduction (p.12) Bertrand Russell explains
”If an atomic fact is analyzed as fully as possible (theoretical, not practical
possibility is meant) the constituents finally reached may be called simples
or objects.”
”(…) the naming of simples is shown to be what is logically first in logic.”
Therefore in Very basic set theory the naming of individuals is logically first.
2. What is a set?
At an early age you develop the mental picture of ‘what is a set’. You belong to a family or a clan, to the inhabitants of a village or a particular region. Then experience shows there are things that constitute a set. The members of a set are the constituents of that set. Therefore set theory describes
a world of individuals,
with a definite membership relation between sets,
the basic properties of a set are the constituents of that set and the sets that set is a member of,
the only constituent of an individual is that individual itself.
Experience shows a set ‘is equal to’ or ‘is not equal to’ another set. Sets with different constituents are different sets, and different sets have different constituents.
Set theory follows rules of logic but the membership relation is definite. That means the logic of ‘true’ and ‘false’ should be changed to the logic of ‘definiteness’.
In the next section we prove logical formulas preserve definiteness.
3. Logic and definiteness
A ‘formula’ is a description of properties of sets. Thus $s\in t$ is a formula. Formulas are denoted by $\Phi$, $\Psi$, $\ldots$ and are built up from formulas by means of negation $\neg\Phi$ (not $\Phi$) and conjunction $\Phi\wedge\Psi$ ($\Phi$ and $\Psi$). Negation takes precedence over conjunction. The parenthesis (((( and )))) preclude ambiguity and improve legibility. Then $s\notin t$ is short for the formula $\neg(s\in t)$.
If it is not the case that both $\Phi$ and $\neg\Phi$, that is formula $\neg(\Phi\wedge\neg\Phi)$, then ‘$\Phi$ is definite’. For every $s$ and for every $t$ it is not the case that both $s\in t$ and $\neg(s\in t)$ thus formula $s\in t$ is definite.
If $\Phi$ is a formula then $(\Phi)$, $\forall s(\Phi)$ and $\exists s(\Phi)$ are formulas. $\forall s(\Phi)$ means ‘for every set it is the case $\Phi$’. The negation of $\forall s(\neg\Phi)$ is denoted by $\exists s(\Phi)$ thus $\exists s(\Phi)$ means ‘exists a set such that $\Phi$’. The parenthesis following $\forall s$ and $\exists s$ are mandatory. However, without ambiguity $\forall s\exists t(s\in t)$ is short for the formula $\forall s(\exists t(s\in t))$.
Set $s$ in $\forall s$ or in $\exists s$ is not a ‘variable’. $\forall s$ means ‘for every set’, thus also for a set denoted by $s$. And $\exists s$ means ‘exists a set’ denoted by $s$. Set theory is a language.
If there is $s$ in $\Phi$ not guided in $\Phi$ by $\forall s$ or by $\exists s$ then ‘$s$ is free in $\Phi$’. Thus $s$ is free and $t$ is not free in formula $\forall t(s\in t)$.
If $\Phi$ is definite then $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite otherwise exists $s$ such that $(\forall s(\Phi)\wedge\neg\forall s(\Phi))$ and $(\exists s(\Phi)\wedge\neg\exists s(\Phi))$ thus $(\Phi\wedge\neg\Phi)$.
We define the logical connectives $\vee$, $\rightarrow$ and $\leftrightarrow$.
$\Phi\vee\Psi$ ($\Phi$ or $\Psi$ or both) is short for the formula ¬($\neg(\neg\Phi\wedge\neg\Psi)$,
$\Phi\rightarrow\Psi$ ($\Phi$ implies $\Psi$) is short for the formula ($(\neg\Phi\vee\Psi)$,
$\Phi\leftrightarrow\Psi$ ($\Phi$ if and only if $\Psi$) is short for the formula ($(\Phi\rightarrow\Psi)\wedge(\Psi\rightarrow\Phi))$.
There are many rewite rules: ¬($\neg(\neg\Phi)$ is logically equivalent to $\Phi$, $\neg(\Phi\leftrightarrow\Psi)$ is logically equivalent to $(\Phi\leftrightarrow\neg\Psi)$ and $((\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi))$ is logically equivalent to $((\Phi\wedge\Psi)\wedge(\neg\Phi\vee\neg\Psi))$. More rewrite rules in (Mendelson [2015] p.15).
If $\Phi$ is definite then it is not the case that both $\Phi$ and $\neg\Phi$. Thus it is not the case that both ¬($\neg(\neg\Phi)$ and $\neg\Phi$ therefore $\neg\Phi$ is definite.
And $(\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi)$ is logically equivalent to $((\Phi\wedge\neg\Phi)\wedge\Psi)\vee(\Phi\wedge(\Psi\wedge\neg\Psi))$ therefore if $\Phi$ and $\Psi$ are definite then $(\Phi\wedge\Psi)$ is definite.
The logical ‘or’, ‘implies’ and ‘if and only if’ are expressed in terms of logical ‘not’ and logical ‘and’. Thus if $\Phi$ and $\Psi$ are definite then
$\neg\Phi$, $\Phi\wedge\Psi$, $\Phi\vee\Psi$, $\Phi\rightarrow\Psi$, $\Phi\leftrightarrow\Psi$, $(\Phi)$, $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite.
Therefore the building up of formulas preserves definiteness.
4. Russell’s paradox
We prove formula $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$, that is Russell’s paradox
”No set can contain all and only those sets which do not contain themselves”.
Assume exists set $s$ such that for every set $u$ $(u\in s\leftrightarrow u\notin u)$.
The formula applies to every $u$ thus we can choose $s$ therefore
exists $s$ such that $(s\in s\leftrightarrow s\notin s)$.
Formula $(s\in s\leftrightarrow s\notin s)$ is short for $(s\in s\vee s\in s)\wedge(s\notin s\vee s\notin s)$ thus
$(s\in s\wedge s\notin s)$.
Formula $s\in s$ is definite thus it is not the case that $(s\in s\wedge s\notin s)$.
Therefore set $s$ does not exists.
Formula $\exists s\forall u(u\in s\leftrightarrow u\notin u)$ is definite therefore $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$. $\Box$
Then Russell’s paradox $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$ is logically equivalent to
$\forall s\exists u(u\in s\leftrightarrow u\in u)$ thus $\forall s\exists u((u\in s\wedge u\in u)\vee(u\notin u\wedge u\notin u))$.
Therefore the logic of definiteness proves that
for every set $s$, an individual is a constituent of $s$ or a set that is not an individual
is not a constituent of $s$.
5. Axiom ‘specification’
If $\Phi$ is definite and if exists set $u\in s$ such that $\Phi$ then it is not the case that for every set $u$ $(u\in s\rightarrow\neg\Phi)$. Thus we can postulate a set $v$ whose constituents are all and only those constituents of $s$ such that for every constituent of $v$ it is the case $\Phi$. That is axiom specification (also known as ‘separation’).
  • if $u$ is free in $\Phi$ and $\Phi$ is definite then
    $\forall s(\exists u(u\in s\wedge\Phi)\rightarrow\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge\Phi)))$.
Thus the conjunction of mutually exclusive properties does not constitute a set.
We prove: If an individual is a constituent of $s$ then exists $v$ whose constituents
are all and only those individuals that are constituent of $s$.
Formula $(u\in s\wedge u\in u)$ is definite and is not the conjunction of mutually exclusive properties. Apply ‘specification’ to set $s$ to find set $v$
$\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge u\in u))$. Then an individual is a constituent of $v$.
The formula applies to every $u$ thus we can choose $v$
$\exists v(v\in v\leftrightarrow(v\in s\wedge v\in v))$ therefore
$\exists v((v\notin v\vee(v\in s\wedge v\in v))\wedge((v\notin s\vee v\notin v)\vee v\in v))$.
Then $\exists v(v\notin v\vee v\in s)$. Thus $v$ is not an individual, or $v$ is an individual and a constituent of $s$. A set of all individuals is not excluded. $\Box$
We prove: The set of all sets is nonexisting.
If $\exists u(u\in s\wedge u\notin u)$ then formula $(u\in s\wedge u\notin u)$ is definite and is not the conjuction of mutually exclusive properties. Apply ‘specification’ to set $s$ to find set $v$
$\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge u\notin u))$. Then a set is a constituent of $v$.
The formula applies to every $u$ thus we can choose $v$
$\exists v(v\in v\leftrightarrow(v\in s\wedge v\notin v))$ therefore
$\exists v((v\notin v\vee(v\in s\wedge v\notin v))\wedge((v\notin s\vee v\in v)\vee v\in v))$.
Then $\exists v(v\notin v\wedge v\notin s)$. Thus $v$ is not an individual and not a constituent of $s$. Therefore the set of all sets is nonexisting. $\Box$
6. Axioms for ‘equality’
Definition 1.
Set $s$ ‘is equal to’ set $t$, denoted by $s=t$ if $\forall u(u\in s\leftrightarrow u\in t)$.
Set $s$ ‘is not equal to’ set $t$, denoted by $s\neq t$ if $\neg(s=t)$. $\Box$
The axioms for equality are
  • (equality) $\forall s\forall t(s=t\rightarrow\forall u(s\in u\leftrightarrow t\in u))$,
  • (individuals) $\forall s(s\in s\rightarrow\forall u(u\neq s\rightarrow u\notin s))$.
Thus if $s=t$ then $s$ and $t$ have the same constituents and $s$ and $t$ are members of the same sets. And the only constituent of an individual is that individual itself.
Then $\forall s\forall t((s\in s\wedge t\notin t)\rightarrow(s\neq t\wedge t\notin s))$.
7. Axiom ‘pairs’
Any pair of sets constitute a set. We postulate ‘pairs’.
  • (pairs) $\forall s\forall t\exists v\forall u(u\in v\leftrightarrow(u=s\vee u=t))$.
Set $v$ is denoted by $\{s,t\}$ thus $\{s,t\}=\{t,s\}$. The ‘singleton’ $\{s\}$ of $s$ is $\{s,s\}$.
$s\notin s\rightarrow\{s\}\neq s$ otherwise $s\in s$. Therefore $s\notin s\rightarrow\{s\}\notin\{s\}$,
$s\in s\rightarrow\{s\}=s$, the singleton of an individual is that individual itself,
$s\neq t\rightarrow\{s,t\}\notin\{s,t\}$. Thus a pair of different sets is not an individual.
However, if $s\neq t$ and $(s\in t\wedge t\in s)$ then $s\notin s$ and $t\notin t$. Then exists set $\{s,t\}$ thus there is the ‘vicious membership circle’ $(s\in t\wedge t\in s)$.
Therefore we postulate ‘regularity’.
8. Axiom ‘regularity’
The naming of individuals is logically first. Thus every set has a constituent, and vicious membership circles do not exist. Therefore we postulate ‘regularity’.
  • (regularity) $\forall s\exists v(v\in s\wedge\forall u((u\in s\wedge u\in v)\rightarrow u\in u))$.
If $s$ is an individual or a set of individuals then $v$ is an individual. Otherwise exists $v$ and $v$ is not an individual and $s$ and $v$ only have individuals in common.
And it is not the case that exists $s$ such that
$\forall v(v\in s\rightarrow\exists u((u\in s\wedge u\in v)\wedge u\notin u))$.
Therefore any sequence of sets $s$, $t$, $u$, $v$, … , finite (a vicious membership circle) or not finite, such that each term is a member of the previous one
$t\in s$, $u\in t$, $v\in u$ is nonexisting.
Thus $\forall s\forall t(s\neq t\rightarrow(s\notin t\vee t\notin s))$.
9. Axiom ‘union’
The members of the members of a set constitute a set. We postulate ‘union’.
  • (union) $\forall s\exists v\forall u(u\in v\leftrightarrow\exists t(t\in s\wedge u\in t))$.
Set $v$ is denoted by $\bigcup s$. If $s$ is an individual or a set of individuals then $\bigcup s=s$. The ‘union $s\cup t$’ is defined by $s\cup t=\bigcup\{s,t\}$. Then $\bigcup\{\{s\},\{t\}\}=\{s,t\}$ and $s\cup(t\cup u)=(s\cup t)\cup u$. If $s$ and $t$ are individuals then $s\cup t=\{s,t\}$.
We prove: Any sequence of different sets $s$, $\bigcup s$, $\bigcup(\bigcup s)$, … is finite.
Then $s\neq\bigcup s\rightarrow\exists t(t\in s\wedge t\notin t)$ otherwise $s=\bigcup s$.
$\bigcup s\neq\bigcup(\bigcup s)\rightarrow\exists u\exists t(u\in t\wedge t\in s\wedge u\notin u\wedge t\notin t)$ otherwise $\bigcup s=\bigcup(\bigcup s)$,
$\bigcup(\bigcup s)\neq\bigcup(\bigcup(\bigcup s))\rightarrow\exists v\exists u\exists t(v\in u\wedge u\in t\wedge t\in s\wedge v\notin v\wedge u\notin u\wedge t\notin t)$,
Then there is an ‘infinite’ sequence of sets $s$, $t$, $u$, $v$, … such that each term is
a member of the previous one $t\in s$, $u\in t$, $v\in u$, … contradictory to regularity.
Therefore the sequence terminates thus exists $w$ belonging to the sequence and $\bigcup w=w$. If $z=s\cup\bigcup s\cup\bigcup(\bigcup s)\cup\ldots$ $w$ then $z$ is a set and $\bigcup z=z$. Set $z$ is called ‘the transitive closure’ of $s$. For every set the transitive closure exists. $\Box$
The definition of equality and axiom pairs are only meaningful if exist sets the naming of which are logically first. Thus with individuals and axiom regularity. And regularity is used with axiom union to prove that for every set the transitive closure exists. The proof needs no other axioms and no ‘natural numbers’.
10. Natural numbers
In (Halmos [1974] p.47) the Peano Axioms for the set of all natural numbers ”used to be considered as the fountainhead of all mathematical knowledge”. In this section we define ‘number sequences’ satisfying the Peano Axioms for natural numbers.
Definition 2.
Set $s$ is a ‘subset’ of $t$, denoted by $s\subseteq t$, if $\forall u(u\in s\rightarrow u\in t)$.
$s$ is ‘transitive’ if $\forall u(u\in s\rightarrow u\subseteq s)$, or equivalently $\bigcup s\subseteq s$. $\Box$
Every individual and every set of individuals is transitive. If $s$ is transitive then $\bigcup s$ and $s\cup\{s\}$ are transitive.
Theorem 1.
If $s$ is transitive and $\exists u(u\in s\wedge u\notin u)$ then a set of individuals is a
member of $s$: $\exists v(v\in s\wedge v\notin v\wedge\forall u(u\in v\rightarrow u\in u))$.
Proof. Apply regularity to $s$ to find $v\in s\wedge v\notin v$ such that
$\forall u((u\in s\wedge u\in v)\rightarrow u\in u)$. Set $s$ is transitive and $v\in s$ thus $\forall u(u\in v\rightarrow u\in s)$.
Therefore $\exists v(v\in s\wedge v\notin v\wedge\forall u(u\in v\rightarrow u\in u))$. $\Box$
If the only individuals that are constituent of transitive set $s$ are $u$ and $v$ then $\{u,v\}$ is the only possible set of individuals that is a member of $s$ therefore $\{u,v\}\in s\cup\{s\}$.
If $s$ and $t$ are transitive sets with transitive members and the only individuals that are members of $s$ or of $t$ are $u$ and $v$, thus $\forall w((w\in s\cup t\wedge w\in w)\rightarrow w\in\{u,v\})$, then the ‘law of trichotomy’ for $s$ and $t$ is easily proved and reads
$(s\notin s\wedge t\notin t)\rightarrow(s\in t\vee s=t\vee t\in s).$
If $s$ is a transitive set with transitive members then $\bigcup s$ is a transitive set with transitive members and $(s\notin s\rightarrow\bigcup s\notin\bigcup s)$. And $s\notin\bigcup s$ otherwise $s\in s$. Therefore $\bigcup s=s$ or $s=\bigcup s\cup\{\bigcup s\}$.
It is straightforward to define a number sequence satisfying the Peano Axioms (PA) for natural numbers. The first axiom of PA reads: There is a first natural number. Natural numbers are also known as ‘finite ordinals’.
Definition 3.
Set $s$ is a ‘natural number with first number $\alpha$’ if $s$ is a transitive
set with transitive members, $\alpha$ is a pair of individuals and
$\bigcup\alpha=\alpha\wedge s\notin\alpha\wedge\forall u(u\in s\cup\{s\}\rightarrow(u\in\alpha\cup\{\alpha\}\vee\bigcup u\in u)).$
The ‘successor’ of natural number $s$ is $s\cup\{s\}$. $\Box$
Then $\alpha$ is a natural number. In physical reality a first number is called ‘the unit of measurement’. The classical Greek concept of number (arithmos) is ”a finite plurality composed of units, where a unit is whatever counts as one thing in the number under consideration.” (Mayberry [2000] p.21).
If $z$ (‘zero’), $x$ and $y$ are individuals then there is number sequence $S$ with first number $\{z,x\}$ (‘x coordinate’) and there is number sequence $T$ with first number $\{z,y\}$ (‘y coordinate’). All elements of both sequences have $z$ in common.
If $u$ belongs to the $S$ sequence then $x\in u$ implies $u$ is an $x$ coordinate, if $v$ belongs to the $T$ sequence then $y\in v$ implies $v$ is an $y$ coordinate. Then $\{u,v\}$ is an ‘ordered pair’ with an intrinsic intuitive meaning. The Kuratowski ordered pair does not have any intrinsic intuitive meaning. It is ”just a convenient way” to define ordered pairs (Mendelson [2015] p.235, Kanamori [2003] §5).
Any pair of individuals is a first number of a natural number sequence. Therefore ‘the set of all natural numbers’ is nonexisting. However, to postulate ‘set $\omega$ of all natural numbers with first number $\alpha$’ is not excluded. Then $\bigcup\omega=\omega$. And if $\bigcup\beta=\beta$ we can define ‘transfinite ordinal numbers with first number $\beta$’ analogous to natural numbers (finite ordinal numbers) with first number $\alpha$.
In the Introduction (Wittgenstein [1999] p.21) Bertrand Russell writes
”No logic can be considered adequate until it has been shown to be capable of dealing with transfinite numbers”.
Thus the logic in Very basic set theory is adequate.
Conclusion
Very basic set theory is a set theory with individuals, the logic of definiteness is adequate and the naming of individuals is logically first. But individuals are ‘not needed for mathematical purposes’ (Oliver and Smliley [2006] §5.1). Thus Russell’s paradox is not the result of the logic of definiteness otherwise individuals exist, and the naming of a special set is logically first (an ‘empty set’, or an ‘infinite set’). Individuals are banished from ‘mainstream set theory’.
There are more differences with ‘mainstream set theory’. For example, the axiom ‘powerset’ is as usual but Cantor’s Theorem does not hold.
  • (powerset) $\forall s\exists v\forall u(u\in v\leftrightarrow u\subseteq s)$.
Set $v$ is denoted by ${\mathcal{P}}s$. Then ${\mathcal{P}}\{s\}=\{\{s\}\}$ thus the ‘number of members’ of $\{s\}$ is equal to the ‘number of members’ of ${\mathcal{P}}\{s\}$.
Cantor’s Theorem (Zermelo [1908] p.276) reads: For every set $s$,
the ‘number of members’ of $s$ is less than the ‘number of members’ of ${\mathcal{P}}s$.
Thus Cantor’s Theorem does not hold.