### v7

# Very Basic Set theory

Abstract.
The naive conception of set is described as ”Any predicate has an extension”. Then Russell’s paradox is used to prove the inconsistency of naive set theory. If the only constituent of a set is that set itself we call that set an ‘individual’. We show Russell’s paradox permits sets that are individuals thus that is not an inconsistency. Zermelo’s axiomatization also permits individuals. Very basic set theory is a set theory with individuals and has a philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus.
Introduction
Georg Cantor defined a set as (Boolos [1971] p.215)
  • ”a ‘many, which can be thought of as one, i.e. a totality of definite
    elements that can be combined into a whole by a law’.”
That means a set is ‘determined’ by its elements. Sets are denoted by $s$, $t$, $u$, …. Symbol $\in$ is the ‘membership relation’ symbol. Then $s\in t$ means $s$ is a ‘constituent’ (also called ‘an element’ or ‘a member’) of the totality $t$. The negation of $s\in t$ is denoted by $s\notin t$. Set $s$ is an ‘individual’ if the only constituent of $s$ is $s$ itself. A set with only $s$ and $t$ as constituents is denoted by $\{s,t\}$ (Zermelo [1908] p.263). Often the terminology is: ‘set $\{s,t\}$ contains only $s$ and $t$’. But the totality $\{s,t\}$ is ‘thought of as one’ thus a set is a mental picture in your mind. Therefore we prefer to write ‘constituent of a set’, not ‘is contained in a set’.
In (Oliver and Smiley [2006] $\S$5.1) Alex Oliver and Timothy Smiley explain why
”numerous textbooks” prohibit individuals: individuals are not needed for
mathematical purposes.
In (Kanamori [2003] pp.285-288) Akihiro Kanamori describes that with the ‘empty
extension’ and the ‘axiom of restriction’ there is no need for individuals (also
known as atoms or urelemente). Individuals ”are objects distinct from the
null set yet having no members and capable of belonging to sets”.
In (Boolos [1971] pp.216-217) George Boolos notes ”individuals (= non-sets)”.
In section 4.6.5 (Mendelson [2015]) Elliott Mendelson explains
”The words ‘individual’ and ‘atom’ are sometimes used as synonyms for
‘urelement’. Thus, urelements have no members”. And in a footnote (p.304)
”Zermelo’s 1908 axiomatization permitted urelements”.
In (Zermelo [1908] pp.262-265) Ernst Zermelo describes in ”10. Theorem.”
”Für jedes Element $x$ von $M$ ist es definit, ob $x$ $\varepsilon$ $x$ oder nicht; diese Möglichkeit
$x$ $\varepsilon$ $x$ ist an und für sich durch unsere Axiome nicht ausgeschlossen.”
Thus ‘$x$ $\varepsilon$ $x$ or not’ is definite; the possibility ”$x$ $\varepsilon$ xx” is not excluded by our axioms.
Therefore Zermelo’s 1908 axiomatization permits sets ‘containing’ themselves.
The membership relation is definite, that means
  • for every $s$ and for every $t$ applies either $s\in t$ or $s\notin t$, but not both.
Then for every set $s$ it is the case $s\in s$ or $s\notin s$, but not both. Therefore $s\in s$ is possible. But a set that ‘contains’ itself is considered problematic (Boolos [1971] p.217, pp.219-220)
”perhaps the mind ought to boggle at the idea of something’s containing itself.”
”It is important to realize how odd the idea of something’s containing itself is.”
”Here are some things. Now we bind them up into a whole. Now we have a set.”
However, ‘binding up some things’ or ‘containing something’ is not the definition of a set. A set does not ‘contain’ something, the constituents of a set can be thought of as one totality.
Russell’s paradox reads (Boolos [1971] p.217)
”No set can contain all and only those sets which do not contain themselves.”
The negation of Russell’s paradox reads
  • for every $s$ exists a set $u$ such that (($u\in s$ and $u\in u$) or ($u\notin s$ and $u\notin u$)).
If we choose for set $u$ set $s$ then the formula reads ‘for every $s$ ($s\in s$ or $s\notin s$)’. It is the case that ‘for every $s$ ($s\in s$)’ is definite. Therefore Russell’s paradox permits sets ‘containing’ themselves.
In $\S$1 we describe sets and individuals. Atomic facts and elementary propositions are defined which gives the foundation of set theory in the Tractatus (Wittgenstein [1999]). In $\S$2 we prove the formulas of set theory preserve definiteness. In $\S$3 we define equality for sets. Then axioms specification, pairs and union are formulated. We prove regularity and some other important theorems. In $\S$7 natural numbers are defined.
1. What are individuals and sets?
At an early age you develop the mental picture of ‘what is a set’. You belong to a family or a clan, to the inhabitants of a village or a particular region. Then experience shows there are things that constitute a set. The members of a set are the constituents of that set.
Set theory is a language that follows rules of logic to communicate the mental picture of a set with the physical world. Then someone reading, seeing or listening that language forms a mental picture of a set. Set theory is made up of ‘formulas’. A formula is a proposition or predicate about the facts of sets. For every set $s$ and for every set $t$ the fact is the definiteness of $s\in t$. A formula which is not complex or not a combination of other formulas is an ‘atomic fact’. Formula $s\in t$ is an atomic fact, formula $t\in s$ is an atomic fact. Formula $s\notin t$ is the negation of $s\in t$ thus $s\notin t$ is a complex formula therefore $s\notin t$ is not an atomic fact. Atomic facts are what makes formulas true or false.
Definition.
Set $s$ is an individual if the only constituent of set $s$ is set $s$ itself.
A formula presupposes propositions or predicates while propositions or predicates presupposes atomic facts. Then if an atomic fact is analyzed as fully as possible the constituents finally reached are individuals. Individuals cannot be analyzed further. In this way the naming of individuals is what is logically first in logic (Wittgenstein [1999] prop. 3.26) and Bertrand Russell (Introduction p.12).
If $s$ is a set then $s$ is not a proposition or predicate. If $s$ is an individual then $s\in s$ therefore $s$ is the ’propositional connexion’ (Wittgenstein [1999] prop. 4.221), (Anscombe [1963] p.29-31). Thus individuals are ‘elementary propositions’. Therefore in Very basic set theory the naming of individuals is logically first.
2. Logic and definiteness
In this section we prove logical formulas preserve definiteness. Formulas are denoted by $\Phi$, $\Psi$, $\ldots$ and are built up from formulas by means of negation $\neg\Phi$ (not $\Phi$) and conjunction $\Phi\wedge\Psi$ ($\Phi$ and $\Psi$). Negation takes precedence over conjunction. If $\Phi$ is a formula then $(\Phi)$ is a formula. The parenthesis (( and )) preclude ambiguity and improve legibility. Then $s\notin t$ is the formula $\neg(s\in t)$.
If $\Phi$ is a formula then $\forall s(\Phi)$ and $\exists s(\Phi)$ are formulas. $\forall s(\Phi)$ means ‘for every set it is the case $\Phi$’. The negation of $\forall s(\neg\Phi)$ is denoted by $\exists s(\Phi)$ thus $\exists s(\Phi)$ means ‘exists a set such that $\Phi$’. The parenthesis following $\forall s$ and $\exists s$ are mandatory. However, e.g. without ambiguity $\forall s\exists t(s\in t)$ is short for the formula $\forall s(\exists t(s\in t))$.
If there is $s$ in $\Phi$ not guided in $\Phi$ by $\forall s$ or by $\exists s$ then ‘$s$ is free in $\Phi$’. Thus $s$ is free and $t$ is not free in formula $\forall t(s\in t)$.
Set $s$ in $\forall s$ or in $\exists s$ is not a ‘variable’. $\forall s$ means ‘for every set’, thus also for a set denoted by $s$. And $\exists s$ means ‘exists a set’ denoted by $s$. Set theory is a language.
We define the logical connectives $\vee$, $\rightarrow$ and $\leftrightarrow$.
$\Phi\vee\Psi$ ($\Phi$ or $\Psi$ or both) is short for the formula $\neg(\neg\Phi\wedge\neg\Psi)$,
$\Phi\rightarrow\Psi$ ($\Phi$ implies $\Psi$) is short for the formula $(\neg\Phi\vee\Psi)$,
$\Phi\leftrightarrow\Psi$ ($\Phi$ if and only if $\Psi$) is short for the formula ($(\Phi\rightarrow\Psi)\wedge(\Psi\rightarrow\Phi))$.
There are many rewrite rules: $\neg(\neg\Phi)$ is logically equivalent to $\Phi$, $\neg(\Phi\leftrightarrow\Psi)$ is logically equivalent to $(\Phi\leftrightarrow\neg\Psi)$ and $((\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi))$ is logically equivalent to $((\Phi\wedge\Psi)\wedge(\neg\Phi\vee\neg\Psi))$. More rewrite rules in (Mendelson [2015] p.15).
Definition.
Formula $\Phi$ is ‘definite’ if it is not the case that both $\Phi$ and $\neg\Phi$. $\Box$
Theorem.
The building up of formulas preserve definiteness.
Proof. For every $s$ and for every $t$ it is not the case that both $s\in t$ and $\neg(s\in t)$
therefore atomic facts and elementary propositions are definite.
If $\Phi$ is definite then it is not the case that both $\Phi$ and $\neg\Phi$. Thus it is not the
case that both $\neg(\neg\Phi)$ and $\neg\Phi$ therefore $\neg\Phi$ is definite.
If $\Phi$ is definite then $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite otherwise $(\forall s(\Phi)\wedge\neg\forall s(\Phi))$ or
$(\exists s(\Phi)\wedge\neg\exists s(\Phi))$ thus $\exists s(\Phi\wedge\neg\Phi)$.
And $(\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi)$ is logically equivalent to
$((\Phi\wedge\neg\Phi)\wedge\Psi)\vee(\Phi\wedge(\Psi\wedge\neg\Psi))$ therefore if $\Phi$ and $\Psi$ are definite then
$(\Phi\wedge\Psi)$ is definite.
The logical connectives ‘or’, ‘implies’ and ‘if and only if’ are expressed in terms
of logical ‘not’ and logical ‘and’. Thus if $\Phi$ and $\Psi$ are definite then $(\Phi\vee\Psi)$,
$(\Phi\rightarrow\Psi)$ and $(\Phi\leftrightarrow\Psi)$ are definite.
If an atomic fact is analyzed as fully as possible constituents finally reached are
individuals. Atomic facts and elementary propositions are definite.
Therefore the building up of formulas preserve definiteness. $\Box$
3. Equality for sets
Experience shows a set ‘is equal to’ or ‘is not equal to’ another set. A set is determined by its constituents. Thus sets with different constituents are different sets, and the constituents of different sets are different. We define equality for sets.
Definition.
Set $s$ ‘is equal to’ set $t$, denoted by $s=t$ if $\forall u(u\in s\leftrightarrow u\in t)$.
Set $s$ ‘is not equal to’ set $t$, denoted by $s\neq t$ if $\neg(s=t)$. $\Box$
The axioms for equality are
  • (equality) $\forall s\forall t(s=t\rightarrow\forall u(s\in u\leftrightarrow t\in u))$,
  • (individuals) $\forall s(s\in s\rightarrow\forall u(u\neq s\rightarrow u\notin s))$.
Thus if $s=t$ then $s$ and $t$ have the same constituents and $s$ and $t$ are members of the same sets. And the only constituent of an individual is that individual itself.
Then $\forall s\forall t((s\in s\wedge t\notin t)\rightarrow(s\neq t\wedge t\notin s))$.
4. Axiom ‘specification’
If $\Phi$ is definite and if exists set $u\in s$ such that $\Phi$ then it is not the case that for every set $u$ $(u\in s\rightarrow\neg\Phi)$. Thus we can postulate a set $v$ whose constituents are all and only those constituents of $s$ such that for every constituent of $v$ it is the case $\Phi$. That is axiom specification (also known as ‘separation’).
  • if $u$ is free in $\Phi$ and $\Phi$ is definite then
    $\forall s(\exists u(u\in s\wedge\Phi)\rightarrow\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge\Phi)))$.
Thus the conjunction of mutually exclusive properties does not constitute a set.
Theorem.
If an individual is a constituent of set $s$ then exists set $v$ whose
constituents are all and only those individuals that are constituent of $s$.
Proof. Formula $(u\in s\wedge u\in u)$ is definite and is not the conjunction of mutually
exclusive properties. Apply specification to set $s$ to find set $v$
$\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge u\in u))$. Then an individual is a constituent of $v$.
The formula applies to every $u$ thus we can choose $v$
$\exists v(v\in v\leftrightarrow(v\in s\wedge v\in v))$ therefore
$\exists v((v\notin v\vee(v\in s\wedge v\in v))\wedge((v\notin s\vee v\notin v)\vee v\in v))$.
Then $\exists v(v\notin v\vee v\in s)$.
Therefore $v$ is not an individual, or $v$ is an individual and a constituent of $s$. $\Box$
Theorem.
The set of all sets is nonexisting.
Proof. If $\exists u(u\in s\wedge u\notin u)$ then formula $(u\in s\wedge u\notin u)$ is definite and is not the
conjuction of mutually exclusive properties.
Apply specification to set $s$ to find set $v$
$\exists v\forall u(u\in v\leftrightarrow(u\in s\wedge u\notin u))$. Then a set is a constituent of $v$.
The formula applies to every $u$ thus we can choose $v$
$\exists v(v\in v\leftrightarrow(v\in s\wedge v\notin v))$ therefore
$\exists v((v\notin v\vee(v\in s\wedge v\notin v))\wedge((v\notin s\vee v\in v)\vee v\in v))$.
Then $\exists v(v\notin v\wedge v\notin s)$.
Thus $v$ is not an individual and $v$ is not a constituent of $s$.
Therefore the set of all sets is nonexisting. $\Box$
5. Axiom ‘pairs’
Any pair of sets constitute a set. We postulate ‘pairs’.
  • (pairs) $\forall s\forall t\exists v\forall u(u\in v\leftrightarrow(u=s\vee u=t))$.
Set $v$ is denoted by $\{s,t\}$, $s$ and $t$ are the only constituents of $\{s,t\}$.
Then $\{s,t\}=\{t,s\}$ thus $\{s,t\}$ is also called ‘an unordered pair’.
The ‘singleton’ $\{s\}$ of set $s$ is $\{s,s\}$. Then
$s\notin s\rightarrow\{s\}\neq s$ otherwise $s\in s$. Therefore $s\notin s\rightarrow\{s\}\notin\{s\}$,
$s\in s\rightarrow\{s\}=s$, the singleton of an individual is that individual itself,
$s\neq t\rightarrow\{s,t\}\notin\{s,t\}$. Thus a pair of different sets is not an individual.
However, if $s\neq t$ and $(s\in t\wedge t\in s)$ then $s\notin s$ and $t\notin t$. Then exists set $\{s,t\}$ and the ‘vicious membership circle’ $(s\in t\wedge t\in s)$. That is contradictory to ‘the naming of individuals is logically first’. Therefore if $s\neq t$ it is not the case that
$\forall v(v\in\{s,t\}\rightarrow\exists u((u\in\{s,t\}\wedge u\in v)\wedge u\notin u))$.
The naming of individuals is logically first. Thus vicious membership circles do not exist. We prove ‘regularity’.
Theorem.
It is not the case that exists set $w$ such that
$\forall v(v\in w\rightarrow\exists u((u\in w\wedge u\in v)\wedge u\notin u))$.
Proof. A finite sequence of sets $s$, $t$, $u$, …, $z$ such that $s\in t$, $t\in u$, …, $z\in s$
is nonexisting otherwise there are vicious membership circles. And an ‘infinite’
sequence of sets $s$, $t$, $u$, $v$, … such that each term is a member of the previous
one $t\in s$, $u\in t$, $v\in u$ is nonexisting otherwise for every set $x$ belonging to
the sequence exists set $y$ belonging to that sequence such that $y\in x$.
That is contradictory to ‘the naming of individuals is logically first’. $\Box$
Formula $\neg\exists w\forall v(v\in w\rightarrow\exists u((u\in w\wedge u\in v)\wedge u\notin u))$ is definite. Therefore
  • (regularity) $\forall w\exists v(v\in w\wedge\forall u((u\in w\wedge u\in v)\rightarrow u\in u))$.
The formula applies to every $u$ thus we can choose $v$. Then the formula reads
$\forall w\exists v(v\in w\wedge((v\notin w\vee v\notin v)\vee v\in v))$, that is $\forall w\exists v(v\in w)$.
Therefore an ‘empty set’ is nonexisting. If an individual is a constituent of $w$ then $v$ is an individual. Otherwise $v$ is not an individual and $w$ and $v$ only have individuals in common.
6. Axiom ‘union’
The members of the members of a set constitute a set. We postulate ‘union’.
  • (union) $\forall s\exists v\forall u(u\in v\leftrightarrow\exists t(t\in s\wedge u\in t))$.
Set $v$ is denoted by $\bigcup s$. If $s$ is an individual or a set of individuals then $\bigcup s=s$. The ‘union $s\cup t$’ is defined by $s\cup t=\bigcup\{s,t\}$. Then $\bigcup\{\{s\},\{t\}\}=\{s,t\}$ and $s\cup(t\cup u)=(s\cup t)\cup u$. If $s$ and $t$ are individuals then $s\cup t=\{s,t\}$.
Theorem.
Any sequence of different sets $s$, $\bigcup s$, $\bigcup(\bigcup s)$, … is finite.
Proof. Then $s\neq\bigcup s\rightarrow\exists t(t\in s\wedge t\notin t)$ otherwise $s=\bigcup s$.
$\bigcup s\neq\bigcup(\bigcup s)\rightarrow\exists u\exists t(u\in t\wedge t\in s\wedge u\notin u\wedge t\notin t)$ otherwise $\bigcup s=\bigcup(\bigcup s)$,
$\bigcup(\bigcup s)\neq\bigcup(\bigcup(\bigcup s))\rightarrow\exists v\exists u\exists t(v\in u\wedge u\in t\wedge t\in s\wedge v\notin v\wedge u\notin u\wedge t\notin t)$,
Then there is an ‘infinite’ sequence of sets $s$, $t$, $u$, $v$, … such that each term is
a member of the previous one $t\in s$, $u\in t$, $v\in u$, … contradictory to regularity.
Therefore the sequence terminates. Thus exists set $w$ belonging to the sequence
such that $\bigcup w=w$. If $z=s\cup\bigcup s\cup\bigcup(\bigcup s)\cup\ldots$ $w$ then $z$ is a set and $\bigcup z=z$.
Set $z$ is called ‘the transitive closure’ of $s$. $\Box$
The definitions of equality and axiom pairs are only meaningful if exist sets the naming of which are logically first. Thus with individuals and regularity. And regularity is used with axiom union to prove that for every set the transitive closure exists. The proof needs no other axioms and no ‘set of all natural numbers’.
7. Natural numbers
In (Halmos [1974] p.47) the Peano Axioms for the set of all natural numbers ”used to be considered as the fountainhead of all mathematical knowledge”. In this section we define ‘transitive sets’ and ‘number sequences’ satisfying the Peano Axioms for natural numbers.
Definition.
Set $s$ is a ‘subset’ of $t$, denoted by $s\subseteq t$, if $\forall u(u\in s\rightarrow u\in t)$.
$s$ is transitive if $\forall u(u\in s\rightarrow u\subseteq s)$, or equivalently $\bigcup s\subseteq s$. $\Box$
Every individual and every set of individuals is transitive. If $s$ is transitive then $\bigcup s$ and $s\cup\{s\}$ are transitive.
Theorem.
If $s$ is transitive and $\exists u(u\in s\wedge u\notin u)$ then a set of individuals is a
member of $s$: $\exists v(v\in s\wedge v\notin v\wedge\forall u(u\in v\rightarrow u\in u))$.
Proof. Apply regularity to $s$ to find $v\in s\wedge v\notin v$ such that
$\forall u((u\in s\wedge u\in v)\rightarrow u\in u)$. Set $s$ is transitive and $v\in s$ thus $\forall u(u\in v\rightarrow u\in s)$.
Therefore $\exists v(v\in s\wedge v\notin v\wedge\forall u(u\in v\rightarrow u\in u))$. $\Box$
If the only individuals that are constituent of transitive set $s$ are $u$ and $v$ then $\{u,v\}$ is the only possible set of individuals that is a member of $s$ therefore $\{u,v\}\in s\cup\{s\}$.
If $s$ and $t$ are transitive sets with transitive members and the only individuals that are members of $s$ or of $t$ are $u$ and $v$, thus $\forall w((w\in s\cup t\wedge w\in w)\rightarrow w\in\{u,v\})$, then the ‘law of trichotomy’ for $s$ and $t$ is easily proved and reads
$(s\notin s\wedge t\notin t)\rightarrow(s\in t\vee s=t\vee t\in s).$
If $s$ is a transitive set with transitive members then $\bigcup s$ is a transitive set with transitive members and $(s\notin s\rightarrow\bigcup s\notin\bigcup s)$. And $s\notin\bigcup s$ otherwise $s\in s$. Therefore $\bigcup s=s$ or $s=\bigcup s\cup\{\bigcup s\}$.
It is straightforward to define a number sequence satisfying the Peano Axioms for natural numbers. The first axiom of the Peano Axioms reads ‘there is a first natural number’. Natural numbers are also known as ‘finite ordinals’.
Definition.
Set $s$ is a ‘natural number with first number $\alpha$’ if $s$ is a transitive
set with transitive members, $\alpha$ is a pair of individuals and
$\bigcup\alpha=\alpha\wedge s\notin\alpha\wedge\forall u(u\in s\cup\{s\}\rightarrow(u\in\alpha\cup\{\alpha\}\vee\bigcup u\in u)).$
The ‘successor’ of natural number $s$ is $s\cup\{s\}$. $\Box$
Then $\alpha$ is a first natural number. In physical reality a first number is called ‘the unit of measurement’. The classical Greek concept of number (arithmos) is ”a finite plurality composed of units, where a unit is whatever counts as one thing in the number under consideration.” (Mayberry [2000] p.21).
If $z$ (‘zero’), $x$ and $y$ are individuals then there is number sequence $S$ with first number $\{z,x\}$ (‘x coordinate’) and there is number sequence $T$ with first number $\{z,y\}$ (‘y coordinate’). All elements of both sequences have $z$ in common.
If $u$ belongs to the $S$ sequence then $x\in u$ implies $u$ is an $x$ coordinate, if $v$ belongs to the $T$ sequence then $y\in v$ implies $v$ is an $y$ coordinate. Then $\{u,v\}$ is an ‘ordered pair’ with an intrinsic intuitive meaning. The Kuratowski ordered pair does not have any intrinsic intuitive meaning. It is ”just a convenient way” to define ordered pairs (Mendelson [2015] p.235, Kanamori [2003] §5).
Any pair of individuals is a first number of a natural number sequence. Therefore ‘the set of all natural numbers’ is nonexisting. However, it is not excluded to postulate ‘set $\omega$ of all natural numbers with first number $\alpha$’. Then $\bigcup\omega=\omega$. Therefore we can define ‘transfinite ordinal numbers with first number $\omega$’ analogous to natural numbers with first number $\alpha$. Therefore the logic in Very basic set theory is capable to deal with transfinite numbers.
Conclusion
Very basic set theory is a set theory that has its philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus. It is a set theory with individuals, the logic is capable to deal with transfinite numbers and the naming of individuals is logically first. There is no ‘empty set’ and regularity is not an axiom but a theorem.
There are more differences with mainstream set theory. For example, the axiom ‘powerset’ is as usual but Cantor’s Theorem does not hold.
  • (powerset) $\forall s\exists v\forall u(u\in v\leftrightarrow u\subseteq s)$.
Set $v$ is denoted by ${\mathcal{P}}s$. Then ${\mathcal{P}}\{s\}=\{\{s\}\}$ thus the ‘number of members’ of $\{s\}$
is equal to the ‘number of members’ of ${\mathcal{P}}\{s\}$.
Cantor’s Theorem (Zermelo [1908] p.276) reads
for every set $s$, the ‘number of members’ of $s$ is less than the ‘number of
members’ of ${\mathcal{P}}s$.
Thus Cantor’s Theorem does not hold. $\Box$